OpenStudy (anonymous):

A waiter can find his/her gross pay for the week, with a 15% tip included, by using this formula: Gross Pay = (Hourly Wage)·(Hours Worked) + (0.15)·(Cost of Meals) Suppose that Serena worked 35 hours at $3.60 an hour and served $1210 in meals. How much did she make that week? Answer (A) $1213.60 (B) $1517.50 (C) $289.60 (D) $307.50

OpenStudy (anonymous):

pay=35*3.6+1210*(15/100)

OpenStudy (anonymous):

o_0 Im Lost

OpenStudy (ghazi):

it's 307$

OpenStudy (anonymous):

\[pay=35\times3.6+1210\times15\div100\] solve it use calculator

OpenStudy (ghazi):

307.5$

OpenStudy (anonymous):

Thanks :)

OpenStudy (ganpat):

Gross Pay = (Hourly Wage)·(Hours Worked) + (0.15)·(Cost of Meals) Gross Pay = 3.6(35) + (0.15) (1210) Gross pay = 307.5 $